Topic: Recursion

int puzzle (int start, int end)

{

if (start > end)

return start - end;

else if (start == end)

return start + end;

else

return end * puzzle (start + 1, end - 1)

}

What is the output of the following statement?

cout << puzzle (5, 10) << endl;

The answer is 720. But how and why?
Answer #1

puzzle(5, 10)
= 10 * puzzle(6, 9)
= 10 * 9 * puzzle(7, 8)
= 10 * 9 * 8 * puzzle(8, 7)
= 10 * 9 * 8 * (8 - 7)
= 10 * 9 * 8 * 1
= 10 * 72
= 720
Answer: 720